OpenStudy (anonymous):

Convert the polar equation r 2 = 2 sin 2θ to a Cartesian equation.

OpenStudy (abb0t):

What's \(sin(2 \theta)\) first of all. Remember your trig identities.

OpenStudy (jdoe0001):

I got a answer, but kinda funky :/

OpenStudy (jdoe0001):

$$ r^2= 4cos(\theta)sin(\theta)\\ \text{both sides times }r^2\\ r^4= 4(rcos(\theta))(rsin(\theta))\\ x^4+y^4 = 4xy $$

OpenStudy (jdoe0001):

was thinking about that... that'd just ... give me a longer trinomial hehe

OpenStudy (jdoe0001):

ever used http://mathb.in/ ?

OpenStudy (jdoe0001):

I think so, good ideahehhe

OpenStudy (loser66):

I join math club, OMG. all them are genius, all them are experts , I feel as if I am a dumb. there

OpenStudy (loser66):

however, I feel stable when I know that my higher level is sponsored by that site. hehehhe I posted many problems here, 7/10 got no answer.